OpenStudy (anonymous):

essay: show all work. A gardenter wants to create a triangular garden in the shape of a right trangle with shortest side length x-3y ft. and the middle length side x+6y ft. what is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simploest correct mathematical form, including units. can someone help me with this one please

OpenStudy (anonymous):

 in a right triangle, the shortest and middle length side will be the height and base of the triangle, as shown in the diagram above that c will always be the side with the greatest value thus , area of the garden = 0.5 * ( x-3y) * (x+6y) = 0.5 ( x^2 + 6xy - 3xy -18y^2) = 0.5 (x^2 +3xy + 18y^2) = 0.5x^2 + 1.5xy + 9y^2 ft^2

OpenStudy (anonymous):

so it is written 0.5x^2+1.5xy+9y^2/2

OpenStudy (anonymous):

0.5x^2+1.5xy+9y^2 the final term is not 9y^2/2

OpenStudy (anonymous):

so i just write it 0.5x^2+1.5xy+9y^2

OpenStudy (anonymous):

yeah with the units of ft ^ 2

OpenStudy (anonymous):

k thanks